# Plug-and-Play Transformer Modules for Test-Time Adaptation

### 10 Jan 2024 v2

@@ 5 Experiment
Table 3: Results on Digits. We train the source models using 4 digits datasets to perform inference on the remaining dataset. The table clearly shows that PLUTO outperforms all of the baselines consistently
@@
Single Source-Best 53.5 98.4 84.5 83.0 94.5 82.7
Source-Worst 21.4 42.2 51.6 12.3 17.1 28.9
TENT-Best 56.1 98.4 84.9 87 95.2 84.3
TENT-Worst 17.6 54.2 59.6 11.4 15.5 31.6
Multi TENT-Ens 62.4 97.8 87.6 55.1 78.4 76.2
TENT-Ens-SA 62.8 98.8 88.7 71.2 86.7 81.6
PLUTO (Ours) 63.1 98.9 89.6 86.4 93.7 86.4
Table 4: Diminishing return. On ImageNet-C (severity level=2) dataset, we train 15 modules for 15 domains, along with the module selector $\mathcal{G}$. For the test-time adaptation performance evaluation, we pick 14 of them to adapt to the rest one. $M$ is the number of selected module, $N$ is the size of the set of all modules. We report the averaged accuracy over 15 target domains for different choice of $M$.
Description $M=1$ $M=3$ $M=5$ $M=8$ $M=11$ $M=14=N-1$
Unlabeled update to module selector $\mathcal{G}$, Adaptation with $M$ modules selected 56.9 59.4 63.7 64.1 64.3 64.4
No update to module selector $\mathcal{G}$, Zero-shot select $M$ source modules, No LN tuning adaptation 42.5 47.3 51.3 53.1 53.3 53.5
Table 5: Larger module store. On ImageNet-C dataset, there are 15 different corruptions with 5 different severity levels, leading to at most $5\times 15=75$ different domains. For the test-time adaptation performance evaluation, we pick one target domain and select $M=14$ source modules to adapt to it. We evaluated the module selection performance on ImageNet-C dataset. Note that, given more source domains for pretraining, for larger $N$, the module selector $\mathcal{G}$ is trained on more source domains, which leads to more complexity. We report the averaged accuracy over $N$ target domains for different choice of $N$.
Description $N=15$ $N=30$ $N=50$ $N=75$
Unlabeled update to module selector $\mathcal{G}$, Adaptation with $M=14$ modules selected 64.4 65.5 65.8 66.2
No update to module selector $\mathcal{G}$, Zero-shot select $M=14$ source modules, No LN tuning adaptation 53.5 55.4 56.5 57.2
Table 6: Few-shot performance of PLUTO. On ImageNet-C (severity level=2) dataset, we train 15 modules for 15 domains. For the test-time adaptation performance evaluation, we pick 14 of them to adapt to the rest one. We randomly pick $U<B$ samples in the test batch to update module selector, make module selection, and LN tuning adaptation. We report the averaged accuracy over 15 target domains for different choice of $U$. $\mathcal{G}$ is pretrained on 15 source domains in a supervised manner.
Description $U=B=128$ $U=64,B=128$ $U=32,B=128$ $U=16,B=128$ $U=4,B=128$ $U=0$, Zero-shot
Baseline Fully-finetune on $T_{target}$ 76.7 $N/A$ $N/A$ $N/A$ $N/A$ $N/A$
Supervised PET on $T_{target}$ 72.1 $N/A$ $N/A$ $N/A$ $N/A$ $N/A$
Randomly select $M=5$ modules, No adaptation $N/A$ $N/A$ $N/A$ $N/A$ $N/A$ 42.2
PLUTO Supervised update to module selector $\mathcal{G}$ on $T_{target}$, Adaptation with $N=14$ modules 68.2 67.5 67.1 65.3 63.1 $N/A$
Unlabeled update to module selector $\mathcal{G}$, Adaptation with $M=N=14$ modules selected 64.4 63.1 62.2 60.9 57.4 50.1
Unlabeled update to module selector $\mathcal{G}$, Adaptation with M=5 modules selected 63.7 62.2 61.2 59.4 56.3 47.3
Table 7: Comparing the issue of forgetting to the baselines on CIFAR-10, this table shows that PLUTO displays significantly slower forgetting rates over an extended cycle of repetitive test distributions in comparison to the baseline methods. It’s worth noting that the "online" methods, which rely on additional domain knowledge, serve as upper bounds for their respective continuous adaptation counterparts.
Time $t$
Round 1 3 5 All
Conditions Clean Snow Fog Frost Clean Snow Fog Frost Clean Snow Fog Frost Avg
TENT-Best-Online 89.9 87.7 87.6 90.4 89.9 87.7 87.6 90.4 89.9 87.7 87.6 90.4 88.9
TENT-Ens-Online 94.1 91.1 93.8 92.5 94.1 91.1 93.8 92.5 94.1 91.1 93.8 92.5 92.9
TENT-Ens-SA-Online 94.3 91.4 94 93.1 94.3 91.4 94 93.1 94.3 91.4 94 93.1 93.2
PLUTO-Online 95.1 93.3 94.4 94.4 95.1 93.3 94.4 94.4 95.1 93.3 94.4 94.4 94.3
TENT-Best 89.9 78.9 82.1 70.2 53.7 45.8 50.6 47 40.7 29.7 31.1 27.4 53.9
TENT-Ens 94.1 85.2 83.1 77.5 61.8 54.4 52.3 47.5 35.2 30.4 30.8 28.7 56.7
TENT-Ens-SA 94.3 88.2 83.7 80.6 66.7 57.5 59.7 62.8 57.4 40.6 30.8 43.1 63.7
PLUTO 95.1 92.5 93.9 92.2 89.1 92.6 89.5 88.6 87.3 87.8 86.9 87.9 90.3
In the main paper, we employ parameter-efficient fine-tuning (PET) in the form of VPT [23, 13] as our module to demonstrate the effectiveness of our approach. We will supplement our method’s application to other PET modules in the supplementary material to illustrate the generality of our approach. The effectiveness of our approach is evaluated on both stationary and continuous test time adaptation benchmarks.During the testing phase, we plan to evaluate our approach under two scenarios: (1) Test batches are sampled from a stationary distribution (2) Test batches are sampled from a evolving distribution.
  • Office-Home [50] The Office-Home dataset consists of four domains: Art (Ar), Clipart (Cl), Product (Pr), and Real World (Re), each containing 65 classes. For evaluating our method in the test-time adaptation scenario, we selected three of these domains as source domains and used the remaining domain as the test data. We conducted online inference on the test data in batches using the source domains, while the test data had its labels removed.
  • ImageNet-C. [17] ImageNet-C dataset consists of 15 diverse corruption types applied to validation images of ImageNet[6]. It consists algorithmically generated corruptions from noise, blur, weather, and digital categories with 5 severity levels. We regard different corruptions with different severity level as different domains. Thus we will have $5\times 15=75$ domains at most.
  • Digits-Five. [36] consists of five-digit datasets: MNIST (mt), MNIST-M (mm), USPS (up), SVHN (sv), Synthetic Digits (sy). There are 10 classes corresponding to digits ranging from 0 to 9. We employed four of the domains as sources and reserved the remaining domain for testing.
In the context of continuous test-time adaptation, we employ the following benchmarks.
  • CIFAR-10C. [17] The CIFAR-10[26] dataset is an image classification dataset consisting of 50,000 training images and 10,000 test images. Based on CIFAR-10, CIFAR-10C [17], introduced 15 different types of noise at different severity levels (1 $\rightarrow$ 5) to the 10,000 test images. This dataset has since been utilized in various studies related to continuous test-time adaptation.
5.1 Baseline Methods
Our main baseline is TENT[52], which is designed for single-source test-time adaptation. We also explore some variants of TENT. To assess adaptation performance, we adopt a setup similar to that in [1]. We apply each source model to the test data independently, resulting in TENT-Best and TENT-Worst, representing the best and worst adaptation performances among these models, respectively. The best/worst results of directly applying source models on target test data without adaptation are noted as Source-Best and Source-Worst. Additionally, we calculate the average pre-softmax prediction (ensemble) over all sources and refer to it as TENT-Ens. To make fair comparison, we also have sharpness-aware (SA) version for the above baselines.
For adaptation on stationary distributions, as our method is based on the ensemble of pre-trained ViT models with pre-trained modules on unlabeled data, we will compare our approach with the following baseline methods: (1) Full model fine-tuning of a pretrained ViT model. (2) Freezing most parameters of a pretrained ViT model and supervising training a module $\theta_{T}$ on the target domain. This comparison will help evaluate the effectiveness of our method against these different fine-tuning/training approaches.
For adaptation on continuous test-time adaptation, to emphasize the issue of forgetting, we contrast our approach with the online variants of the mentioned methods. In the online scenario, we have access to an oracle that signals domain changes, presenting an ideal situation. Here, the source models are reset whenever the oracle identifies a new domain. However, it just serves as an upper-bound solely for continuous adaptation scenarios where domain knowledge is accessible, which is not a practical real-world scenario.
5.2 Implementation details
We use ViT-Base-16[8] model pretrained on JFT-300M dataset for all our experiments. For all the experiments we use a target batch size of $B=128$, as used by TENT [52]. For more details on implementation and experimental setting see supplementary.
5.3 Object recognition and digit classification
We report the results of object recognition on Office-home [50] and Digits-Five [36] dataset in Table. 23. We compute the accuracy for each incoming test batch and subsequently present the results by averaging the accuracy values across all the batches. For Office-home dataset, the number of source modules is $N=3$, and we use all $M=N$ (in Alg. 1) to do the inference on the target domain. Similar setup for Digits-Five dataset, except $M=N=4$ here.
5.4 Module selection
In the practical application scenarios of our method, there may be a considerable number of candidate pretrained modules corresponding to various source domains. However, not all of these source domains have a significant impact on the adaptation to the target domain. Therefore, as illustrated in Algorithm  1, out of all the $N$ pretrained source modules, we only utilize $M(<N)$ modules for domain adaptation. The core idea of our approach is: (1) to identify the most relevant source domains/source modules to the target domain and (2) to assign greater weights to these modules in the model ensemble to provide better-weighted pre-softmax logits. We will demonstrate that we have achieved these two points through experiments from two aspects:
(1) Diminishing return. We select a growing number of modules from a fixed set: We believe that by measuring the average weight allocated to each source by the module selector on a target test batch, we can reflect the general importance of these source domains to this batch’s relationships. Therefore, by selecting the top $M(M<N)$ source modules corresponding to the highest average weights and using only these M sources for adaptation.
(2) Larger module store Adding more pretrained modules (leading to larger $N$) to the current module set while keeping the number of modules involved in adaptation the same. PLUTO can still identify the most important modules while the size of the module store increases.
5.5 Few-shot adaptation
As shown in Alg. 1, our approach involves training an atTENTion-style module selector responsible for generating weights to ensemble the outputs of various source models. This process is designed to tailor predictions for each target sample. By doing so, PLUTO can effectively account for sample-specific preferences in ensembling different source models, particularly when provided with limited labeled or unlabeled target data, as shown in Table. 6
We showcase the few-shot performance via demostrating results on ImageNet-C. On ImageNet-C dataset, we train 15 modules for 15 corruption domains (with severity level=2). Then for $t$-th target batch $D_{T}^{(t)}=\{x^{(t)}_{i}\}_{i=1}^{B}$ Use we pick different $U(<B)$ samples to update module selector $\mathcal{G}$ and make selection in Alg. 1.
5.6 Forgetting issue
We extensively experiment with this dataset to assess our model’s performance in the face of dynamic test distributions. We introduce three types of corruption noise from CIFAR-10C: Snow, Fog, and Frost, each with a severity level of 5. These noises are added to the original CIFAR-10 training set, leading to the training of three source modules, each dedicated to one of these noise types. Additionally, we train a module on the clean training set of CIFAR-10. In other words, we trained $N=4$ modules on noisy CIFAR-10C dataset. Evaluation of forgetting conducted on other datasets will be included in the supplementary material.
During testing, we evaluate these models sequentially on test data corresponding to each noise type. We progressively increase the severity level until reaching the maximum, followed by a gradual decrease in severity level in reverse order. The sequence structure is as follows: (Snow) severity: $\cdots\rightarrow 1\rightarrow 2\rightarrow 3\rightarrow 4\rightarrow 5\rightarrow 4\rightarrow 3\rightarrow 2\rightarrow 1\rightarrow 2\rightarrow\cdots$ (Fog) severity:$\cdots\rightarrow 1\rightarrow 2\rightarrow 3\rightarrow 4\rightarrow 5\rightarrow 4\rightarrow 3\rightarrow 2\rightarrow 1\rightarrow 2\rightarrow\cdots$, and so forth.
We repeat this entire cycle five times to observe how each model’s performance deteriorates over time due to forgetting. In each of the five cycles, the test data remain identical, allowing for a fair comparison with previous cycles. We present the results of this experiment in Table.7, distinguishing between the online and continual versions with a horizontal line. The online version represents the best-case scenario that relies on additional oracle domain information.
6 Conclusion
We present PLUTO, a Test Time Adaptation (TTA) algorithm for transformers, combining multiple source parameter-efficient tuning (PET) modules during test time. PLUTO achieves test accuracy exceeding the best individual source model and prevents catastrophic forgetting. Experiments reveal two key findings: (i) PLUTO outperforms TENT-Best, TENT-Ens, and TENT-Ens-SA for various test distributions and achieves good few-shot performance. (ii) Over repeated cycles, baseline methods degrade significantly, while PLUTO maintains high performance with slower forgetting.

### 08 Feb 2024 v3

@@ Appendix A Sharpness-aware domain adaptation
Table A: Results on Office-Home. We train the source modules using 3 datasets to perform inference on the remaining dataset.
Source Method Ar Cl Pr Re Avg
Baseline (In Main paper) TENT-Ens 63.1 49.2 75.6 71.0 64.7
PLUTO 66.0 53.6 79.2 76.3 68.7
Variants of PLUTO Module selection 63.4 50.1 76.1 73.4 65.8
+ SA 65.1 52.9 78.3 75.6 67.9
+ Filter 63.7 51.5 76.8 74.1 66.3
Appendix B Parameter efficiency
Table B: Number of parameters for different models’ scales and their corresponding PET module size in our setting. ViT-T/S/B/L stands for "Tiny, Small, Base, Large", corresponding to different pretrained ViT sizes. The bolded numbers are the size of the PET modules we applied in our experiments. We highlighted the two size of PET modules in our experiments.
ViT-T ViT-S ViT-B ViT-L
Full model 5,543,716 21,704,164 85,875,556 303,404,132
Adapter 58,564 116,932 233,668 417,984
VPT 37,732 75,364 150,628 299,108
Header 19,300 38,500 76,900 102,500
Parameter-efficient tuning (PET) methods align naturally with model ensemble techniques(Lester et al., 2021), particularly in terms of parameter efficiency. In contrast to other models where an ensemble of $N$ models results in $N$ times more model parameters, an ensemble of $N$ distinct PET modules merely entails $N$ times more modules. This is because all models to be ensembled share the same pre-trained model that conditions the PET modules. Therefore, although PLUTO is model ensemble test-time adaptation approach, the additional model parameters introduced by the ensemble are only the PET modules, i.e. $N\times|\theta|$. In our approach, we train an attention-based module selector $\mathcal{G}$ which includes four projection layers and two layer norms. It required $d\times d^{\prime}_{x}+d^{\prime}_{x}\times d^{\prime}+v\times d^{\prime}_{l}+d^{\prime}_{l}\times d^{\prime}+4d^{\prime}$ parameters ($\approx$ 0.8M in our experiment). It only required about less than 1% of a pretrained ViT-base model ($\approx$ 86M, Table. B).
Appendix C Details on different PET methods
C.1 Visual-Prompt Tuning (VPT)
With a pre-trained Transformer (ViT) model as our starting point, we introduce a set of $p$ continuous embeddings in the input space, each of dimension $d$, referred to as "prompts." During fine-tuning, only the prompts specific to the task are updated, while the Transformer backbone remains frozen. We applied VPT-Shallow(Jia et al., 2022) as following:
VPT-Shallow. Prompts are inserted into the first Transformer layer only. Each prompt token is a learnable $d-$ dimensional vector. A module, which is a collection of $p$ (which is the prompt length) prompts is denoted as $\mathbf{P}=\{\mathbf{v}\in\mathbb{R}^{d}|k\in\mathbf{N},1\leq k\leq p\}$, the shallow-prompted ViT is:
$[x_{1},\mathbf{Z}_{1},\mathbf{E}_{1}]={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}L_{1}}([{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}x_{0}},\mathbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}P}},\mathbf{E}_{0}])$ (A)
$[x_{i},\mathbf{Z}_{i},\mathbf{E}_{i}]={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}L_{i}}([x_{i-1},\mathbf{Z}_{i-1},\mathbf{E}_{i-1}]),i=2,\cdots,l$ (B)
$\mathbf{y}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\text{Head}}(x_{l}),$ (C)
where $\mathbf{Z}_{i}\in\mathbb{R}^{p\times d}$ represents the features computed by the $i$-th Transformer layer, and $[x_{i},\mathbf{Z}_{i},\mathbf{E}_{i}]\in\mathbb{R}^{(1+p+P)\times d}$ ($P$ is the number of patches that a 2D image input is divided into). The color above indicate learnable and frozen parameters, respectively. For ViTs, $x_{i}$ is invariant to the location of prompts since they are inserted after positional encoding. The overall parameter count for Adapters in an $l$-layer Transformer can be calculated as $|\theta|=p\times d.$
C.2 Adapter
In the conventional configuration, a transformer model incorporates two Adapters per layer (Baevski & Auli, 2018). Each Adapter layer is composed of $2\times k\times d$ parameters, accounting for both the down and up-projection matrices. Here, $k$ represents the input dimension size, while $d$ refers to the bottleneck dimension of the Adapter. The overall parameter count for Adapters in an $l$-layer Transformer can be calculated as $|\theta|=2\times l\times 2\times k\times d.$
C.3 Results for different PET modules
Following the setup in main paper Sec. 4.3, we replace VPT with Adapters as the PET module, and demonstrate the stationary distribution adaptation results in Table. C and Table. D. In can be concluded that our method, PLUTO, remains good performance across different PET modules.
Table C: Results on Office-Home. We train the source modules using 3 datasets to perform inference on the remaining dataset ($M=N=3$). The difference here is that we applied Adapter instead of VPT in the main paper.
Source Method Ar Cl Pr Re Avg
Single TENT-Best 65.6 52 77.9 72.1 66.9
TENT-Worst 47.5 47.1 57.6 65.3 54.4
CoTTA-Best 65.3 50.9 75.8 70.9 65.7
CoTTA-Worst 46.6 46.2 56.1 63.8 53.2
EaTA-Best 65.5 51.8 76.6 72.0 66.5
EaTA-Worst 46.7 46.3 55.2 61.9 52.5
Multi TENT-Ens 63.1 49.2 76.6 72.9 65.5
CoTTA-Ens 65.4 49.9 77.2 71.8 66.1
EaTA-Ens 65.2 50.7 76.3 71.6 66.0
PLUTO (Ours) 65.7 55.4 79.1 77.1 69.3
Table D: Results on Digits. We train the source modules using 4 digits datasets to perform inference on the remaining dataset ($M=N=4$). The difference here is that we applied Adapter instead of VPT in the main paper.
@@ C.3 Results for different PET modules
Single TENT-Best 54.5 97.6 83.7 86.6 94.5 83.4
TENT-Worst 16.4 54.0 59.2 11.7 15.7 31.4
CoTTA-Best 53.5 98.5 85.8 84.8 94.5 83.4
CoTTA-Worst 16.2 53.9 58.6 10.3 16.3 31.1
EaTA-Best 55.4 97.4 83.6 85.9 92.8 83.0
EaTA-Worst 16.9 51.7 57.9 8.7 16.5 30.3
Multi TENT-Ens 61.8 98.8 85.8 53.2 77.1 75.3
CoTTA-Ens 56.2 98.6 85.5 87.0 95.3 84.5
EaTA-Ens 56.3 96.7 83.8 86.4 94.5 83.5
PLUTO (Ours) 64.0 99.1 89.9 87.3 95.5 87.2
Appendix D Hyperparameters
The experimental setup of the attention-based module selector $\mathcal{G}$ is listed as in Table. E. Table. F summarizes the optimization configurations we used. Implementation details for each tuning method on source/target tasks are also included.
Table E: Hyperparameters of the module selector $\mathcal{G}$ for different datasets.
Model Hyper-param CIFAR100-C ImageNet-C Office-Home 5-Digits
ViT-Base $d^{\prime}_{x}$ 64 128 32 32
$d^{\prime}_{l}$ 64 128 32 32
$d^{\prime}$ 64 128 64 32
Dropout % 0 0 0 50
Table F: Hyperparameters and implementation details for each tuning methods.
Full, Adapter VPT
Optimizer AdamW SGD
Optimizer momentum N/A 0.9
$base\_lr$ search range {0.001, 0.0001, 0.0005, 0.005} {50., 25., 10., 5., 2.5, 1.,0.5, 0.25, 0.1, 0.05}
Weight decay range {0.01, 0.001, 0.0001, 0.0}
LR schedule cosine decay
Warm up epochs 10
Total epochs 100
Appendix E Details of dataset corruptions
We summarize these corruptions types by example in Fig. A. The order of these corruptions is same as the order in Table. 6 and Table. 7
Figure A: Examples of each corruption type in the image corruptions benchmark. While synthetic, this set of corruptions aims to represent natural factors of variation like noise, blur, weather, and digital imaging effects.
Appendix F Theoretical Insights
PLUTO’s goal is to determine the ideal weights $\{w(x)^{j}\}_{j=1}^{N}$ for every source, combining them to create the target predictor. We will demonstrate that, given logical assumptions about the source and target distributions, there is a straightforward selection of the target predictor. This selection can match or surpass the performance of the best source model when applied directly to the target data.
[…]
$\theta_{S}^{k}=\text{argmin}_{\theta}\mathcal{L}(Q_{S}^{k},\theta),\forall 1\leq k\leq N,$
then the target predictor can be experssed as
$\theta_{T}(x)=\sum_{k=1}^{N}\frac{\lambda_{k}Q_{S}^{k}(x)}{\sum_{j=1}^{N}\lambda_{j}Q_{S}^{j}(x)}\theta^{k}_{S}(x).$
By applying Lemma 1,2 in (Ahmed et al., 2021), we can know that the pseudo-label loss (Eqn. 2) and the supervised loss are both less than or equal to the loss induced by the best source predictor, i.e.
$\mathcal{L}(Q_{T},\theta_{T})\leq\min_{1\leq j\leq N}\mathcal{L}(Q_{T},\theta_{S}^{j}).$
Which is consistent with the results shown in Table. 23.